1.75 inch dia hole in a .75 inch plate of A36 steel. A force of 34000 lbs vertical to plane is applied to the hole. Use bearing stress to determine stress placed vertical to the plane and at 45 degrees to the plane
Answer #1

SOLUTION:

Projected area for bearing stress : A = 1.75*0.75 = 1.3125 in2

Stress in the vertical plane = Sv = P/A = 34000/1.3125

Sv = 25904.76 psi : ANSWER

Now Sv is the principal stress in vertical direction.

Now stress in the plane 45 degree to the vertical plane = S = (Sv/2) + (Sv/2) cos( 2*45)

S = 25904.76/2 = 12952.38 psi

S = 12952.38 psi : ANSWER

Please rate.